# Aeos: The Impact of Pop III Initial Mass Function and Star-by-Star Models in Galaxy Simulations

### v1

@@ 2 Methods
This study uses the numerical methods detailed in Brauer et al. (2025). We focus on a comparative analysis between the fiducial Aeos cosmological simulation and additional simulations: (1) an Aeos simulation with an altered Pop III IMF with a higher characteristic mass $M_{\rm char}$ and upper mass cutoff, and (2) seven comparison simulations that omit the full star-by-star feedback and enrichment models, instead using single stellar population particles and IMF averaged metal yields. The comparison simulations are identical except for stochastic differences to show how stochasticity affects results. Below, we provide a concise summary of the important methods from Brauer et al. (2025) and elaborate on the modifications pertinent to the additional simulations. The differences in the simulations are summarized in Table 1.
@@ 2.1 Similarities Across Simulations
Star formation is modeled stochastically in cold, dense gas regions with $n>10^{4}$ cm-3 that exhibit converging flows having $\nabla\cdot\mathbf{v}<0$, assuming a star formation efficiency per free-fall time of $e_{\rm ff}=2\%$. The star formation algorithm distinguishes Pop II from Pop III stars using a threshold of total gas metallicity $Z>10^{-5}Z_{\odot}$ (Ji et al., 2014; Chiaki et al., 2015; Schneider et al., 2012; Tumlinson, 2006). Pop III star formation in lower metallicity gas further requires a molecular hydrogen fraction $f_{\rm H_{2}}>0.005$, consistent with high-resolution simulations (Susa et al., 2014; Kulkarni et al., 2021).
@@
Above $M_{\text{char}}$, it behaves as a Salpeter IMF, and below $M_{\text{char}}$ it is exponentially suppressed (Chabrier, 2003). We also include an upper mass limit $M_{\rm max}$. The Pop II IMF is sampled from a Kroupa (2001) distribution with a mass range of 0.08–120 M.
@@
All simulations employ a modified version of grackle (Smith et al., 2017) to handle non-equilibrium chemistry involving nine species (H, H+, He, He+, He++, e-, H-, H2, and H+${}_{2}^{+}$) and associated radiative cooling and heating processes. The chemical network includes dust-mediated H2 formation, a UV background that extends to high redshift and is scaled to be continuous with Haardt & Madau (2012) at $z=10$, and a Lyman-Werner background model from Emerick et al. (2019) and Wise et al. (2012b) that adopts updated rates at high redshift from Qin et al. (2020).
@@ 2.2 Differences Between Simulations
In the Aeos simulations, Aeos10 and Aeos20, star particles represent individual stars sampled from their respective IMFs except for Pop II stars below 2 M, which are aggregated into single particles due to their negligible feedback and enrichment contributions on simulation timescales. Star formation occurs from gas reservoirs of $\geq 100$ M.
In the comparison simulations, Pop III stars are represented by individual particles, but Pop II stars are represented by star cluster particles with a minimum mass of 1000 M. For Pop II star formation, stars are formed from gas reservoirs of $\geq 1000$ M.
@@ 3.1 Ionization
Figure 4: Stellar masses and halo masses of star-forming halos in the different simulations. (a) Comparing the star-forming halos in Aeos10 vs. Aeos20. Aeos10 has many more small star-forming halos than Aeos20. This is due to higher ionization in Aeos20 (see Figure 1). The distribution of halo masses for all halos in all simulations are nearly identical, as expected.
@@
The higher ionization in Aeos20 suppresses the formation of small galaxies. Previous work has found that ionization-driven photoheating of the gas in and around dwarf galaxies suppresses star formation (Efstathiou, 1992; Bullock et al., 2000), particularly in halos with $M_{\text{halo}}\lesssim 10^{9}$ M (e.g., Thoul & Weinberg, 1996; Okamoto et al., 2008; Dawoodbhoy et al., 2018). In our simulations, all galaxies exist in small halos and we find they are highly sensitive to the rate of reionization. As seen in Figure 4, Aeos10 has far more galaxies at low halo masses. For $M_{*}<100$ M, Aeos10 has 60 star-forming halos while Aeos20 has only 14 at $z=14.5$. Meanwhile, the distributions of halo masses for all halos are essentially identical for Aeos10, Aeos20, and the comparison simulations, as expected for simulations with the same dark matter initialization. Another mechanism, Lyman-Werner (LW) radiation, can also suppress early star formation by dissociating H2, the primary coolant in metal-poor gas (e.g., Safranek-Shrader et al., 2012). Both Aeos10 and Aeos20 have identical LW backgrounds, however, so the suppression of small galaxy formation in Aeos20 vs. Aeos10 is isolated to the differences in ionizing photons.
@@ 3.2 Chemical Enrichment
When we vary the Pop III IMF, we thus produce different chemical abundances in the subsequent Pop II stars. In Figure 5, we show how the differences in yields translate into the chemical abundances of the oldest Pop II stars. We show the median stellar chemical abundances of first-generation Pop II stars from each galaxy in the simulations that has begun Pop II star formation (17 galaxies in Aeos10, 15 galaxies in Aeos20). We limit the analysis to low-mass stars, as more massive stars would not survive to the present-day.
@@
Pop II stars in Aeos20 exhibit higher [X/Fe] values for light elements such as nitrogen (N) and oxygen (O), reflecting the increased nucleosynthetic yields from more massive Pop III stars due to stronger $\alpha$-capture reactions and rotational mixing (Heger & Woosley, 2010). Carbon (C) is also enhanced in Aeos20, primarily due to the significant production of carbon during helium burning in more massive Pop III stars. In these simulations we do not include pair-instability SNe (Rakavy & Shaviv, 1967; Fraley, 1968; Woosley et al., 2002), which may occur for high-mass ($\gtrsim 140$ M) Pop III stars (e.g. Schneider et al., 2004; Takahashi et al., 2016; de Bennassuti et al., 2017), but this would further enhance the amount of carbon due to their efficient ejection of carbon-rich material. For $\alpha$-elements such as magnesium (Mg), silicon (Si), and calcium (Ca), Aeos20 also tends to show higher [X/Fe] ratios, consistent with the Heger & Woosley (2010) yields and the expectation that more massive Pop III stars are more efficient producers of $\alpha$-elements during their evolution and subsequent SN explosions.
@@
The enhanced carbon yield of more massive Pop III stars results in different amounts of carbon-enhanced metal-poor (CEMP) stars. CEMP stars, identified by their high carbon-to-iron ratios (taken either at [C/Fe] > 0.7 or [C/Fe] > 1.0), provide key insights into the formation of the first stars and galaxies (Beers & Christlieb, 2005; Jeon et al., 2017). Figure 6 shows the cumulative fractions of our simulated CEMP stars as a function of metallicity. Because more massive Pop III stars produce a higher ratio of carbon to iron, Aeos20 has a higher fraction of CEMP stars than Aeos10. While they reproduce the qualitative trends, compared with the observed fractions of CEMP stars in the Milky Way (Placco et al., 2014; Yoon et al., 2018), neither Aeos10 nor Aeos20 overproduce CEMP stars at a given [Fe/H]. Generally, at the lowest metallicities that are most sensitive to Pop III enrichment, Aeos10 stars have lower fractions than observations and Aeos20 stars have higher fractions. This shows that the differences in Pop III initial mass functions between Aeos10 and Aeos20 significantly affect CEMP fractions and therefore points toward an improved IMF that can better reproduce Milky Way observations.
@@
In observations, the very faintest dwarf galaxies appear to exhibit a metallicity floor rather than further extending the mass-metallicity relationship found for the more luminous galaxies (e.g., Ahvazi et al., 2024; Heiger et al., 2024). Across three orders of magnitude in luminosity ($10^{2}-10^{5}$ M), ultra-faint dwarf galaxies show a nearly constant mean metallicity, scattered around [Fe/H]$\sim$-2.6 (Fu et al., 2023). This contrasts with the expectation that less massive galaxies should have lower average metallicities. The observed plateau is not yet fully understood, but may be caused by Pop III enrichment (Wheeler et al., 2019), a ceiling on mass outflows (Kravtsov & Manwadkar, 2022), or the possibility that these galaxies were once more massive and later lost stellar mass due to tidal disruption (e.g., Kirby et al., 2013). In the Aeos simulations, we find that Pop III enrichment results in a median metallicity floor of approximately Z = $10^{-4}$ Z (see also Figure 8 of Brauer et al. 2025) with increasing stellar metallicities for galaxies of $\sim$105 M and greater. This is a slightly lower metallicity floor than seen in observed dwarf galaxies, implying either additional Pop III enrichment in observed galaxies or that an additional mechanism is necessary.
@@ 4 Comparison to Simulations Without Individual Stellar Feedback
Figure 9: Comparing the star-forming halos in Aeos10 with those in a typical comparison simulation. The stellar mass distributions differ; the comparison simulations generally form more galaxies around $\sim 1000$ M. This is because in the comparison simulations, the Pop II stars form in particles of 1000 M, biasing any halo that has begun Pop II star formation to contain at least $1000$ M. The halo mass distributions are more similar because the dark matter modeling is the same between simulations.
@@
The cumulative Pop III and Pop II star formation of the comparison simulations can be seen in the top panel of Figure 8. While differences exist in the overall star formation history, the differences are similar in magnitude to stochastic differences seen between identically initialized simulations, with the exception that Pop II formation tends to begin slightly earlier in the comparison simulations. However, the difference in onset of Pop II star formation is not large, and could be due to stochastic differences. As seen in the bottom of Figure 8, Aeos10 and the comparison simulations do not significantly differ in their overall star formation history. Aeos10 and Aeos20 also resemble each other, though Aeos20 starts star formation slightly behind Aeos10. In general, the bulk properties of total star formation and stellar mass-halo mass relation are not highly affected by the differences between the Aeos simulations and the comparison simulations.
Figure 9 shows the mass distribution of star-forming halos in Aeos10 vs. a comparison simulation. As expected, when looking only at halo mass, the distributions are similar; the halo mass is dominated by dark matter, which does not differ between the simulations. The stellar masses, however, differ. The comparison simulations have a tendency to form galaxies of at least a few thousand solar masses because the Pop II star particles form in masses of $\sim$1000 M. The number of Pop II halos in each comparison simulation differs from just a few to a few dozen, but in aggregate, there is a trend towards more galaxies around 1000 M than in the Aeos simulations. The Aeos galaxies, on the other hand, can include Pop II galaxies with stellar masses as low as a few tens of solar masses (see Brauer et al., 2025). Additionally, both the Aeos simulations and comparison simulations can have stripped stars, resulting in the peak at a few M.

### v2

@@ 2 Methods
This study uses the numerical methods detailed in Brauer et al. (2025). We focus on a comparative analysis between the fiducial Aeos cosmological simulation and additional simulations: (1) an Aeos simulation with an altered Pop III IMF with a higher characteristic mass $M_{\rm char}$ and upper mass cutoff, and (2) seven comparison simulations that omit the full star-by-star feedback and enrichment models, instead using single stellar population particles and IMF averaged metal yields. The comparison simulations are identical except for differences arising primarily from stochastic IMF sampling during star formation. We run seven comparison simulations to show how stochasticity affects results. Below, we provide a concise summary of the important methods from Brauer et al. (2025) and elaborate on the modifications pertinent to the additional simulations. The differences in the simulations are summarized in Table 1.
@@ 2.1 Similarities Across Simulations
Star formation is modeled stochastically in cold, dense gas regions with $n>10^{4}$ cm-3 that exhibit converging flows having $\nabla\cdot\mathbf{v}<0$, assuming a star formation efficiency per free-fall time of $e_{\rm ff}=2\%$. The star formation algorithm distinguishes Pop II from Pop III stars using a threshold of total gas metallicity $Z>10^{-5}Z_{\odot}$ (Ji et al., 2014; Chiaki et al., 2015; Schneider et al., 2012; Tumlinson, 2006). Pop III star formation in lower metallicity gas further requires a molecular hydrogen fraction $f_{\rm H_{2}}>0.0005$, consistent with high-resolution simulations (Susa et al., 2014; Kulkarni et al., 2021).
@@
Above $M_{\text{char}}$, it behaves as a Salpeter IMF, and below $M_{\text{char}}$ it is exponentially suppressed (Chabrier, 2003). We also include an upper mass limit $M_{\rm max}$. In all simulations, every Pop III star is represented by a single star particle. The Pop II IMF is sampled from a Kroupa (2001) distribution with a mass range of 0.08–120 M.
@@
[…]
@@ 2.2 Differences Between Simulations
In the Aeos simulations, Aeos10 and Aeos20, star particles represent individual stars sampled from their respective IMFs except for Pop II stars below 2 M, which are aggregated into single particles due to their negligible feedback and enrichment contributions on simulation timescales. Star formation occurs from gas reservoirs of $\geq 100$ M. Individual star particles are formed by sampling from the IMF until the gas reservoir is depleted, at which time star formation ends until more gas meets the conditions for star formation.
In the comparison simulations, Pop III stars are represented by individual particles, but Pop II stars are represented by star cluster particles with a minimum mass of 1000 M. For Pop II star formation, stars are formed from gas reservoirs of $\geq 1000$ M. A single star particle is formed with a stellar mass distribution determined by sampling from the IMF.
@@ 3.1 Ionization
Figure 4: Stellar masses (left) and halo masses (middle) of star-forming halos in the different simulations. Aeos10 has many more small star-forming halos than Aeos20. This is due to higher ionization in Aeos20 (see Figure 1). The distribution of halo masses for all halos (right) in all simulations are nearly identical, as expected.
@@
The higher ionization in Aeos20 suppresses the formation of small galaxies. Previous work has found that ionization-driven photoheating of the gas in and around dwarf galaxies suppresses star formation (Efstathiou, 1992; Bullock et al., 2000), particularly in halos with $M_{\text{halo}}\lesssim 10^{9}$ M (e.g., Thoul & Weinberg, 1996; Okamoto et al., 2008; Dawoodbhoy et al., 2018). In our simulations, all galaxies exist in small halos and we find they are highly sensitive to the rate of reionization. As seen in Figure 4, Aeos10 has far more galaxies at low halo masses. For $M_{*}<100$ M, Aeos10 has 60 star-forming halos while Aeos20 has only 14 at $z=14.5$. Meanwhile, the distributions of halo masses for all halos are essentially identical for Aeos10, Aeos20, and the comparison simulations, as expected for simulations with the same dark matter initialization. Another mechanism, Lyman-Werner (LW) radiation, can also suppress early star formation by dissociating H2, the primary coolant in metal-poor gas (e.g., Safranek-Shrader et al., 2012). Both Aeos10 and Aeos20 have identical LW backgrounds, however, so the suppression of small galaxy formation in Aeos20 vs. Aeos10 is isolated to the differences in ionizing photons. At the beginning of the simulation ($\sim 120$ Myr), Aeos10 briefly surpasses Aeos20 in ionizing photons (see Figures 1 and 3). This is due to stochastic differences during the beginning of Pop III star formation, when there are very few stars, and a slightly delayed onset of Pop III star formation in Aeos20. An ensemble of simulations with Aeos10 and Aeos20 parameters would be preferred to better understand stochastic differences, but is currently unfeasible due to the computational cost of the simulations.
@@ 3.2 Chemical Enrichment
When we vary the Pop III IMF, we thus produce different chemical abundances in the subsequent Pop II stars. In Figure 5, we show how the differences in yields translate into the chemical abundances of the oldest Pop II stars. We show the median stellar chemical abundances of first-generation Pop II stars from each galaxy in the simulations that has begun Pop II star formation (17 galaxies in Aeos10, 15 galaxies in Aeos20). We limit the analysis to low-mass stars, e.g., the $\leq 0.8$ M stars that would survive to present day in the unresolved star particles.
@@
Pop II stars in Aeos20 exhibit higher [X/Fe] values for light elements such as nitrogen (N) and oxygen (O), reflecting the increased nucleosynthetic yields from more massive Pop III stars due to stronger $\alpha$-capture reactions and rotational mixing (Heger & Woosley, 2010). Carbon (C) is also enhanced in Aeos20, primarily due to the significant production of carbon during helium burning in more massive Pop III stars. In these simulations we do not include pair-instability SNe (Rakavy & Shaviv, 1967; Fraley, 1968; Woosley et al., 2002), which may occur for high-mass ($\gtrsim 140$ M) Pop III stars (e.g. Schneider et al., 2004; Takahashi et al., 2016; de Bennassuti et al., 2017), but this would further enhance the amount of carbon due to their efficient ejection of carbon-rich material. For $\alpha$-elements such as magnesium (Mg) and calcium (Ca), Aeos20 also tends to show higher [X/Fe] ratios, consistent with the Heger & Woosley (2010) yields and the expectation that more massive Pop III stars are more efficient producers of $\alpha$-elements during their evolution and subsequent SN explosions.
@@
The enhanced carbon yield of more massive Pop III stars results in different amounts of carbon-enhanced metal-poor (CEMP) stars. CEMP stars, identified by their high carbon-to-iron ratios (taken either at [C/Fe] > 0.7 or [C/Fe] > 1.0), provide key insights into the formation of the first stars and galaxies (Beers & Christlieb, 2005; Jeon et al., 2017). Figure 6 shows the cumulative fractions of our simulated CEMP stars as a function of metallicity. Because more massive Pop III stars produce a higher ratio of carbon to iron, Aeos20 has a higher fraction of CEMP stars than Aeos10. While they reproduce the qualitative trends, compared with the observed fractions of CEMP stars in the Milky Way (Placco et al., 2014; Yoon et al., 2018), neither Aeos10 nor Aeos20 consistently reproduce observations CEMP stars at a given [Fe/H]. Generally, at the lowest metallicities that are most sensitive to Pop III enrichment, Aeos10 stars have lower fractions than observations and Aeos20 stars have higher fractions. This shows that the differences in Pop III initial mass functions between Aeos10 and Aeos20 significantly affect CEMP fractions and therefore points toward an improved IMF that can better reproduce Milky Way observations.
@@
[…]
@@ 4 Comparison to Simulations Without Individual Stellar Feedback
Figure 9: Comparing the star-forming halos in Aeos10 with those in a typical comparison simulation. The stellar mass distributions differ; the comparison simulations generally form more galaxies around $\sim 1000$ M. This is because in the comparison simulations, the Pop II stars form in particles of 1000 M, biasing any halo that has begun Pop II star formation to contain at least 10001000 M. The halo mass distributions are more similar because the dark matter modeling is the same between simulations.
The cumulative Pop III and Pop II star formation of the comparison simulations can be seen in the top panel of Figure 8. While differences exist in the overall star formation history, the differences are similar in magnitude to stochastic differences seen between identically initialized simulations, with the exception that Pop II formation tends to begin slightly earlier in the comparison simulations. However, the difference in onset of Pop II star formation is not large, and could be due to stochastic differences. As seen in the bottom of Figure 8, Aeos10 and the comparison simulations do not significantly differ in their overall star formation history. In general, the bulk properties of total star formation and stellar mass-halo mass relation are not highly affected by the differences between the Aeos simulations and the comparison simulations.
Aeos10 and Aeos20 also resemble each other, though Aeos20 starts star formation slightly behind Aeos10 ($\Delta z\simeq 3$, $\sim 15$ Myr). This arises from the interplay between the IMF and our discrete star particle spawning scheme. Both simulations adopt identical initial conditions and the same Pop III eligibility threshold; however, once a cell first satisfies those conditions we stochastically draw individual stellar masses from the Pop III IMF and instantiate one star particle per draw until the local gas reservoir (initialized when $\gtrsim 100\,M_{\odot}$ of eligible gas is present) would be overdrawn, at which point the event terminates without forming the overdrawn star. Because Aeos20 uses a more top-heavy IMF, it has a substantially higher probability that an early draw exceeds the available $\sim 100\,M_{\odot}$ reservoir, potentially aborting the first few prospective events and forcing the halo to accumulate additional gas before a successful Pop III burst occurs.
Figure 9 shows the mass distribution of star-forming halos in Aeos10 vs. a comparison simulation. When looking only at halo mass, the distributions are similar. This appears to be because (1) the halo mass is dominated by dark matter, which does not differ between the simulations, and (2) the handling of Pop III stars is the same in the two simulations. In comparison to Figure 4, the masses of Pop III stars are more influential on the masses of star-forming halos than the feedback and resolution of Pop II stars.
The stellar masses, however, differ. The comparison simulations have a tendency to form galaxies of at least a few thousand solar masses because the Pop II star particles form in masses of $\sim$1000 M. The number of Pop II halos in each comparison simulation differs from just a few to a few dozen, but in aggregate, there is a trend towards more galaxies around 1000 M than in the Aeos simulations. The Aeos galaxies, on the other hand, can include Pop II galaxies with stellar masses as low as a few tens of solar masses (see Brauer et al., 2025). Additionally, both the Aeos simulations and comparison simulations can have stripped stars, resulting in the peak at a few M.